bill533:

If f(x) = -x + 4 and g(x) = x^2 - 5, then g(f(3)) = ? A) -8 B) -4 C) -3 D) 0 E) 3

bill533:

If f(x) = -x + 4 and g(x) = x^2 - 5, then g(f(3)) = ? We first evaluate f(3) = -3 + 4 = 1 So.....now we have g (1) Plug 1 into g and evaluate

bill533:

@dude

bill533:

how to do that

bill533:

@umm

Babyplier:

this isn't that hard

Babyplier:

first, plug 3 in for all the x's in f(x)

Babyplier:

then, you plug whatever you get from that into g(x). and then you get ur answer.

bill533:

We have to do the functions inside, and work our way out. Calculate f(3), and using the outcome, galculate the g(x) function of that. f(3) = -3+4 = 1 g(1) = 1^2-5

Babyplier:

yeah

bill533:

-4

bill533:

Find f(4). f(x) = x2 + 2x - 3 A) 5 B) 11 C) 13 D) 21 E) 27

bill533:

is it D

Babyplier:

sadly, I can't give you a direct answer unless I wanna be banned

Babyplier:

but it pretty easy. just put 4 in for all of the x's

Babyplier:

@bill533